find the surface area and volume of a sphere which has a diameter of 48 cm.: find the surface area and volume of a sphere which has a diameter of 48 cm.
Answer by jim_thompson5910(9927)

Since the diameter is 48, the radius is 48/2=24

Surface Area:
Solved by pluggable solver: To calculate Surface area of a sphere

Geometry Surface Area and Volume of 3 Dimensional Figures
The Formula for Surface area Sphere is
Area=4*pi*radius*radius
Area=4*pi*24*24=7238.2294656


Hence, The Surface Area of Sphere of radius 24 is 7238.2294656.


For more on this topic, See the lessons on Geometry: Area and Surface Area

Some more is on Surface Area and Volume and Perimeter and Area.



Volume:
Solved by pluggable solver: FIND volume of a sphere
Volume of a sphere V=(4/3)*pi*r^3 = (4/3)*pi*24^3 = 57905.8348032